Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places. 95% confidence; n = 360, x = 45
Answer #1

Given that,

n = 360

x = 45

Point estimate = sample proportion = \hat p = x / n = 45/3600.125

1 - \hat p = 1-0.125=0.875

At 95% confidence level

\alpha = 1 - 95%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.125*0.875) /360 )

= 0.0342